At the end of four years, Romeo deposits R500 into an investment account. If the interest rate on the account is 10% per annum compounded yearly, determine the value of his investment at the end of 4 years.
1
Expert's answer
2021-09-28T15:54:19-0400

Here, P=R500, r=0.1,

and t=4

"A=P(1+r)^t"


"=500(1+0.1)^4"


"= 500(1.1)^4"

"=R732.05"

The value of his investment at the end of 4 years is R732.05